# Minimal Construction of Graphs with Maximum Robustness

### v1

@@ Abstract
The notions of network $r$-robustness and $(r,s)$-robustness have been earlier introduced in the literature to achieve resilient control in the presence of misbehaving agents. However, while higher robustness levels provide networks with higher tolerances against the misbehaving agents, they also require dense communication structures, which are not always desirable for systems with limited capabilities and energy capacities. Therefore, this paper studies the fundamental structures behind $r$-robustness and $(r,s)$- robustness properties in two different ways. (a) We first explore and establish the tight necessary conditions on the number of edges for undirected graphs with any nodes must satisfy to achieve maximum $r$- and $(r,s)$-robustness. (b) We then use these conditions to construct two classes of undirected graphs, referred as to $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs), that provably achieve maximum robustness with minimal numbers of edges. We finally validate our work through some sets of simulations.
@@ 1 Introduction
Distributed multi-agent systems have many practical uses, ranging from information gathering [1], target tracking [2], to collaborative decision making [3]. A fundamental multi-agent systems’ objective is consensus, under which multiple agents reach agreement on common state values. However, it is known that performance of consensus deteriorates significantly when one or more misbehaving agents share wrong, or even adversarial, information with the system. Resilient consensus in the presence of misbehaving agents has been studied extensively [4, 5, 6, 7, 8, 9, 10, 11]. In [4, 6, 8], an algorithm called Weighted Mean-Subsequence-Reduced (W-MSR) was introduced to allow the non-misbehaving agents (often called normal) to reach consensus despite the presence of misbehaving agents. Unlike approaches such as those in [12, 13, 14], which rely on trust assessments - often at the cost of increased computational complexity - W-MSR instead achieves resilient consensus through a rule-based filtering strategy that discards extreme values. When applied to graphs satisfying the topological properties of $r$-robustness or $(r,s)$-robustness, W-MSR guarantees that normal agents can converge to a value within the convex hull of their initial conditions, even in the presence of a bounded number of adversaries.
@@
In general, a higher robustness in a communication graph allows normal agents to tolerate a larger number of misbehaving agents while achieving consensus through the W-MSR algorithm [4]. Consequently, maximizing the robustness of a graph is highly desirable. However, achieving higher robustness by definition requires a higher number of edges, often referred to as edge counts. This results in dense and complex graph structures, which can pose challenges for certain systems due to practical constraints such as limited communication range, bandwidth, and energy. This clear tradeoff between robustness levels and edge counts motivates our interest in finding the fundamental structures behind these robust graphs. We first discover the smallest edge counts possible to achieve maximum levels of robustness. Leveraging this, in this paper, we then introduce and construct two classes of graphs that achieve the maximum robustness with a minimal set of edges for any given number of nodes.
Several studies have attempted to indirectly address the tradeoff between robustness and edge counts by relaxing the connectivity of graphs. Some have explored time-varying network topologies to reduce connectivity requirements [15, 16, 17]. In [18], the concept of trusted agents, who are always deemed fault-free, is introduced to maintain robustness with fewer edges. Furthermore, robustness of multi-hop communication graphs has been investigated in [19]. While these works aim to relax the connectivity requirements for $r$- and $(r,s)$-robustness to some extent, their results still remain constrained by the structures fundamental to the robustness definitions. A deeper understanding of the fundamental or minimal graph structures that preserve $r$- and $(r,s)$-robustness would further provide insights into these methods for effectively reducing these connectivity constraints.
@@
Despite its merits, identifying the fundamental structures for given levels of robustness is particularly challenging due to its combinatorial nature. In fact, determining $r$ robustness is a coNP-complete problem [8]. Although algorithms for computing exact robustness exist [20, 21], they scale poorly with the number of nodes. This computational bottleneck makes brute-force search approaches impractical. Although there have been works to speed up computations with sampling-based [22], heuristics [23], or machine learning-based approaches [24], they either provide approximations without guarantees or require specific assumptions about graph topologies. In addition, other works under-bound a given graph’s robustness with its algebraic connectivity $\lambda_{2}$ (the second smallest eigenvalue of a graph’s Laplacian matrix) [25, 26] or with its minimum degree [27, 28], which are generally simpler and faster to compute. However, these bounds can be too loose for many graph structures, limiting its practicality in general applications.
@@
To circumvent the combinatorial issue, many researches have focused on the construction and analysis of specific graph classes with predetermined $r$- and $(r,s)$-robustness properties. A complete graph with $n$ nodes is $\lceil\frac{n}{2}\rceil$-robust, indicating that $\lceil\frac{n}{2}\rceil$ is the maximum level of $r$-robustness any graph of $n$ nodes can achieve [4]. Authors of [29] presented a class of circulant digraphs whose robustness levels are directly related to their connectivity parameter $k$. A preferential-attachment method is investigated in [6, 4, 8] for scalable robust graph constructions. The authors in [30] study the construction of a class of undirected $r$-robust graphs with minimum number of nodes to achieve certain robustness. Furthermore, robustness levels of various graph structures, such as geometric and random graphs [31, 32, 8, 33, 34] have been explored.
@@
Still, understanding the fundamental structural properties of general $r$- and $(r,s)$-robust graphs remains an open problem. While the authors in [35] have established the lower bounds on edge counts and construction mechanisms for undirected $(2,2)$-robust graphs with any number of nodes, these results do not generalize to other robustness levels. In [6, 4], the minimum degree requirements for $r$- and $(r,s)$-robust graphs are presented. Nevertheless, these bounds are local in the sense that they only apply to individual nodes. The concept of minimal edge sets for $(r,s)$-robust graphs has been introduced in [36], though specific scalable construction mechanisms are not discussed.
1.0.1 Contributions
In this paper, we explicitly explore the necessary structural properties of graphs with maximum $r$- and $(r,s)$-robustness, building upon our previous work [37], in which we have constructed graphs that achieve maximum $r$-robustness with a minimal set of edges. In this paper, we aim to paint the complete picture of the tradeoff between the robustness levels and edge counts by also considering $(r,s)$-robustness and presenting related analyses. In addition, we further generalize the class of graphs presented in [37] [^1] In [37], we referred to graphs that achieve maximum robustness with minimal sets of edges as $(2r-1,r)$- and $(2r,r)$-robust graphs. To avoid confusion with the standard definition of $(r,s)$-robustness, we adopt a new naming convention in this paper, detailed later in this section.. Finally, we empirically support our results through simulations by showing that removing even a single edge immediately reduces the networks’ robustness and thus the maximum number of the misbehaving agents that can be tolerated in achieving consensus through the W-MSR algorithm. In summary, we make the following contributions in this work:
@@ 1.0.1 Contributions
    We establish the necessary conditions on the edge counts, in the forms of tight lower bounds, that undirected graphs with any number of nodes must satisfy to achieve the maximum $r$- or $(r,s)$-robustness.
@@
    Based on these necessary conditions, we construct two classes of undirected graphs with a varying number of nodes, referred to as $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs), in the paper, that are proven to achieve maximum $r$- and $(r,s)$-robustness with minimal sets of edges, respectively. Then, we further strengthen the result by proving that the construction mechanisms for the maximum $(r,s)$-robustness are in fact necessary and sufficient.
@@
    Through simulation, we empirically demonstrate that $\gamma$- and $(\gamma,\gamma)$-MERGs maintain maximum robustness in a minimalistic manner, with the removal of any single edge resulting in a decrease in robustness.
The remainder of this paper is organized as follows. Section 2 presents preliminaries and the problem statement. Section 3 and 4 focus on necessary structures for $r$-robustness: Section 3 establishes necessary conditions on edge counts for graphs to achieve maximum $r$-robustness, whereas Section 4 provides systematic constructions of graphs that use the fewest number of edges to achieve maximum $r$-robustness. Similarly, Section 5 and 6 examine the necessary structures for maximum $(r,s)$-robustness, where Section 5 determines the necessary conditions on edge counts for graphs to achieve the maximum $(r,s)$-robustness, and Section 6 discusses how to construct graphs that achieve maximum $(r,s)$-robustness with the smallest edge counts. Section 7 shows the simulation results, and in Section 8 conclusions and discussions of future work directions are given.
@@ 2.1 Notation
We denote a simple, undirected time-invariant graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ and $\mathcal{E}$ are the vertex set and the undirected edge set of graph respectively. We denote the number of nodes $n$. An undirected edge $(i,j)\in\mathcal{E}$ indicates that information is exchanged between nodes $i$ and $j$. The neighbor set of agent $i\in\mathcal{V}$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}\;|\;(i,j)\in\mathcal{E}\}$. The state of agent $i$ at time $t$ is denoted as $x_{i}[t]$. We denote the cardinality of a set $\mathcal{S}$ as $|\mathcal{S}|$. We denote the set of non-negative and positive integers as $\mathbb{Z}_{\geq 0}$ and $\mathbb{Z}_{>0}$.
2.2 Weighted Mean Subsequence Reduced Algoritheorem
@@
  • $w_{ij}[t]\geq\alpha$, $10\leq\alpha<$ 1 $\forall j\in\mathcal{N}_{i}\cup\{i\}$,
@@
Through the protocol given by (1), agents’ states reach consensus as long as the graph has a rooted-out branching (i.e. there exists a node that has paths to all other nodes in the graph) [38, 39]. However, as discussed in [4], this algorithm cannot achieve consensus in the presence of misbehaving agents. In our paper, we focus on two types of misbehaving agents, whose definitions are given below:
Definition 1 (malicious agent)
@@
Definition 2 (Byzantine agent)
@@
Agents that are not misbehaving are called normal agents. There are numerous scope models to describe the number of misbehaving agents in a network [4]. The misbehaving agents are assumed to adopt one model, and we discuss two of them below:
Definition 3 ($\mathbf{F}$-total)
@@
Definition 4 ($\mathbf{F}$-local)
@@
In response to various threat and scope models, algorithms on resilient consensus [4, 15, 40, 5, 12, 13] have become very popular. In particular, the W-MSR (Weighted-Mean Subsequent Reduced) algorithm [4] with the parameter $F$ guarantees normal agents to achieve asymptotic consensus with $F$-total or $F$-local misbehaving agents under certain assumed topological properties of the communication graph. We review these properties and other relevant concepts in the next section.
2.3 $r$- and $(r,s)$-Robustness
Definition 5 ($\mathbf{r}$-reachable)
@@ 2.3 $r$- and $(r,s)$-Robustness
Definition 6 ($\mathbf{r}$-robust)
@@
Definition 7 ($\mathbf{(r,s)}$-reachable)
@@
Definition 8 ($\mathbf{(r,s)}$-robust)
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is $\mathbf{(r,s)}$-robust if for every pair of nonempty, disjoint subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, at least one of the following holds (recall $\mathcal{X}^{r}_{\mathcal{S}_{k}}=\{i\in\mathcal{S}_{k}:|\mathcal{N}_{i}\setminus\mathcal{S}_{k}|\geq r\}$ for $k\in\{1,2\}$):
@@
It can be stated that $(r,s)$-robustness is a generalization of $r$-robustness. If a graph is $(r,s)$-robust, it implies it is also $r$-robust. Conversely, if a graph is $r$-robust, it implies it is at least $(r,1)$-robust. These notions of robustness quantify how many misbehaving agents the network can tolerate while ensuring normal agents in a network to reach consensus on a value within the convex hull of their initial values through the W-MSR algorithm. Specifically, $(2F+1)$-robustness provides a sufficient condition for normal agents to achieve consensus in the presence of $F$-total and $F$-local malicious and Byzantine agents. Similarly, $(F+1,F+1)$-robustness serves as a necessary and sufficient condition for tolerating $F$-total malicious agents [4].

### v2

# arXiv:2507.00415v2[eess.SY] 27 Feb 2026

@@ Abstract
The notions of $r$-robustness and $(r,s)$-robustness of a network have been earlier introduced in the literature to achieve resilient consensus in the presence of misbehaving agents. However, while higher robustness levels enable networks to tolerate a higher number of misbehaving agents, they also require dense communication structures, which are not always desirable for systems with limited communication ranges, energy, and resources. Therefore, this paper studies the fundamental structures behind $r$-robustness and $(r,s)$- robustness properties in two ways. (a) We first establish tight necessary conditions on the number of edges that an undirected graph with an arbitrary number of nodes must have to achieve maximum $r$- and $(r,s)$-robustness. (b) We then use these conditions to construct two classes of undirected graphs, referred as to $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs), that provably achieve maximum robustness with minimal numbers of edges. We demonstrate the effectiveness of our method via comparison against existing robust graph structures and a set of simulations.
@@ 1 Introduction
Distributed multi-agent systems have many practical uses, ranging from information gathering [1], target tracking [2], to collaborative decision making [3]. At the core of many distributed algorithms lies consensus, where multiple agents reach agreement on common state values. However, it is known that performance of consensus deteriorates significantly when one or more misbehaving agents share wrong, or even adversarial, information with the system. As a result, resilient consensus in the presence of misbehaving agents has been studied extensively [4, 5, 6, 7, 8, 9, 10, 11]. In [4, 6, 8], an algorithm called Weighted Mean-Subsequence-Reduced (W-MSR) algorithm was introduced to allow the non-misbehaving agents (often called normal) to reach consensus despite the presence of misbehaving agents. In contrast to reputation- and detection-based approaches [12, 13, 14] that require high computational overhead, the W-MSR algorithm instead ensures resilience through a low-complexity filtering strategy.
The notions of $r$- and $(r,s)$-robustness have been introduced in [4, 6] to characterize the resilience guarantees provided by the W-MSR algorithm. In particular, they quantify how many misbehaving agents the network tolerates while ensuring normal agents in a network to reach consensus within the convex hull of their initial values through the W-MSR algorithm [4]. $(2F+1)$-robustness provides a sufficient condition for normal agents to achieve consensus in the presence of $F$-local Byzantine agents. In contrast, $(F+1,F+1)$-robustness serves as a necessary and sufficient condition for tolerating $F$-total malicious agents [4]. Due to their distinct resilience guarantees, both $r$- and $(r,s)$-robustness have attracted significant attention and have been recently extended to ensure resilience in distributed estimation [15], optimization [16, 17], and learning [18, 19]. Their effectiveness has been demonstrated across a broad range of networked systems, including cooperative robotics and smart grids [20, 21, 22, 17].
@@
Consequently, maximizing the robustness of a graph is highly desirable. However, achieving higher robustness by definition requires a higher number of edges, often referred to as edge counts. This results in dense and complex graph structures, which pose challenges for certain systems due to practical constraints such as limited communication range, bandwidth, and energy. This fundamental tradeoff between robustness and edge counts motivates us to investigate the following question: How can we systematically construct communication graphs that achieve maximum robustness with minimal edge counts?
@@
Despite its clear practical importance, identifying the fundamental structures of robust graphs is particularly challenging due to the combinatorial nature of the problem. In fact, determining $r$-robustness is a coNP-complete problem [8]. Although algorithms for computing exact robustness exist [23, 24], they scale poorly with the number of nodes. This computational bottleneck makes brute-force search approaches impractical. Although there have been works to speed up computations with sampling-based [25], heuristics [26], or machine learning-based approaches [27], they either provide approximations without guarantees or require specific assumptions about graph topologies. In addition, other works under-bound a given graph’s robustness with other graph properties that are generally simpler and faster to compute [4, 22, 28, 20, 29]. However, these bounds are in general too loose for many graph structures.
@@
To circumvent the combinatorial issue, some researches have focused on the construction and analysis of specific graph classes with predetermined $r$- and $(r,s)$-robustness properties. A complete graph with $n$ nodes is $\lceil\frac{n}{2}\rceil$-robust, indicating that $\lceil\frac{n}{2}\rceil$ is the maximum level of $r$-robustness any graph of $n$ nodes achieves [4]. Authors of [30] analyzed robustness of a class of circulant digraphs. A preferential-attachment method is investigated in [6, 4, 8] for scalable robust graph constructions. Furthermore, robustness levels of various graph structures, such as geometric and random graphs [31, 32, 8, 33, 34, 35] have been explored.
@@
Still, understanding the fundamental structural properties of general $r$- and $(r,s)$-robust graphs remains an open problem. While the authors in [36] have established the lower bounds on edge counts and construction mechanisms for undirected $(2,2)$-robust graphs, these results do not generalize to other robustness levels. The authors in [37] study the construction of a class of undirected $r$-robust graphs with the minimum number of nodes. While this is equivalent to constructing graphs with maximum $r$-robustness for a given number of nodes, they do not provide any guarantees on the minimality of the edge set, and their construction method is valid only for specific values of $n$. The concept of minimal edge sets for $(r,s)$-robust graphs has been introduced in [38], though specific scalable construction mechanisms are not discussed.
Instead of investigating the fundamental structures of robust graphs, several studies have sought to relax the connectivity requirements associated with robustness. Some have explored time-varying network topologies to reduce connectivity requirements [39, 40, 41]. In [42], authors introduced trusted agents, who are always deemed fault-free, to achieve robustness with fewer edges. Furthermore, robustness under multi-hop communication schemes has been investigated in [43]. More recently, authors in [44] have proposed relaxing the connectivity requirements by employing detection-based isolation strategies. While these reduce the connectivity requirements associated with $r$- and $(r,s)$-robustness, they only do so by lowering the required robustness level and thus still depend on robust graph structures. This further underscores the need to investigate the fundamental structural properties of robustness.
Contributions
Our work directly investigates the fundamental structures of $r$- and $(r,s)$-robust graphs through two primary ways. First, we derive tight lower bounds on edge counts required for graphs with any number of nodes to achieve maximum robustness. To the best of our knowledge, these bounds are the first to explicitly depend on the level of robustness. Then, for an arbitrary number of nodes, we present two classes of graphs that provably achieve maximum robustness with the minimum number of edges. We refer to these graphs as $\gamma$- and $(\gamma,\gamma)$-Minimal Edge Robust Graphs (MERGs) [^1] In [45], we referred to graphs that achieve maximum $r$-robustness with minimal sets of edges as (2r1,r)(2r-1,r)- and $(2r,r)$-robust graphs. To avoid confusion with the standard definition of $(r,s)$-robustness, we adopt a new naming convention in this paper.. We build upon our previous work [45], in which we only studied $r$-robust graphs with minimal sets of edges, by also considering $(r,s)$-robust graphs. Specifically, we make the following contributions:
@@
    We establish tight lower bounds on the number of edges that an undirected graph with any number of nodes must have to achieve maximum $r$- or $(r,s)$-robustness.
@@
    Based on the lower bounds, we construct two classes of undirected graphs with an arbitrary number of nodes, referred to as $\gamma$- and $(\gamma,\gamma)$-MERGs that are proven to achieve maximum $r$- and $(r,s)$-robustness with minimal edge sets, respectively. Moreover, we prove that the construction mechanisms for $(\gamma,\gamma)$-MERGs are both necessary and sufficient, thereby capturing the entire space of graphs that achieve maximum $(r,s)$-robustness with the fewest edges.
@@
    We evaluate our methods through comparisons with other classes of robust graphs and a series of simulations.
The remainder of this paper is organized as follows. Section 2 presents preliminaries and the problem statement. Section 3 establishes necessary conditions on edge counts for graphs to achieve maximum $r$-robustness, whereas Section 4 provides systematic constructions of $\gamma$-MERGs. Section 5 determines the necessary conditions on edge counts for graphs to achieve maximum $(r,s)$-robustness, and Section 6 presents constructions of $(\gamma,\gamma)$-MERGs. Section 7 compares the proposed constructions with other existing graph constructions. Section 8 presents simulation results, and Section 9 concludes the paper and discusses directions for future research.
@@ 2.1 Notation
We denote a simple, undirected time-invariant graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ and $\mathcal{E}$ are the vertex set and the undirected edge set of graph respectively. We denote the number of nodes $n$. An undirected edge $(i,j)\in\mathcal{E}$ indicates that information is exchanged between nodes $i$ and $j$. The neighbor set of agent $i\in\mathcal{V}$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}\;|\;(i,j)\in\mathcal{E}\}$. The state of agent $i$ at time $t$ is denoted as $x_{i}[t]$. We denote the cardinality of a set $\mathcal{S}$ as $|\mathcal{S}|$. We denote the set of non-negative and positive integers as $\mathbb{Z}_{\geq 0}$ and $\mathbb{Z}_{>0}$. Lastly, we denote by $\lceil\cdot\rceil$ and $\lfloor\cdot\rfloor$ the ceiling and floor operators, respectively.
2.2 Weighted Mean Subsequence Reduced Algorithm
@@
  • $w_{ij}[t]\geq\alpha$, $0\leq\alpha<1$ $\forall j\in\mathcal{N}_{i}\cup\{i\}$,
@@
Through the protocol given by (1), agents are guaranteed to reach consensus on their states as long as the graph is connected [46, 47]. However, consensus may not be reached in the presence of misbehaving agents [4]. In our paper, we assume that the identities of misbehaving agents are unknown, and we focus on two types of misbehaving agents:
Definition 1 (Malicious agent).
@@
Definition 2 (Byzantine agent).
@@
A malicious agent may deviate from the nominal update law but still broadcasts the same value to all of its neighbors. In contrast, a Byzantine agent may both deviate from the protocol and send different values to different neighbors. Hence, a malicious agent is as a special, weaker case of a Byzantine agent. Agents that are not misbehaving are called normal agents. There are numerous models to describe the number of misbehaving agents in a network [4]. We focus on two such models below:
Definition 3 ($\mathbf{F}$-total).
@@
Definition 4 ($\mathbf{F}$-local).
@@
In response to various threat and scope models, algorithms on resilient consensus [4, 39, 48, 5, 12, 13] have become very popular. In particular, the W-MSR (Weighted-Mean Subsequent Reduced) algorithm [4] with the parameter $F$ guarantees normal agents to achieve asymptotic consensus with $F$-total or $F$-local misbehaving agents under certain assumed topological properties of the communication graph.
Definition 5 ($\mathbf{r}$-reachable).
@@
Definition 6 ($\mathbf{r}$-robust).
@@
Definition 7 ($\mathbf{(r,s)}$-reachable).
@@
Definition 8 ($\mathbf{(r,s)}$-robust).
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is $\mathbf{(r,s)}$-robust if for every pair of nonempty, disjoint subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, at least one of the following holds (let $\mathcal{X}^{r}_{\mathcal{S}_{k}}=\{i\in\mathcal{S}_{k}:|\mathcal{N}_{i}\setminus\mathcal{S}_{k}|\geq r\}$ for $k\in\{1,2\}$):
@@
A set $\mathcal{S}$ is $r$-reachable if it contains at least one node with at least $r$ neighbors outside the set, and $(r,s)$-reachable if it contains at least $s$ such nodes. These notions help us define $r$- and $(r,s)$-robustness. A graph is $r$-robust if, for every pair of disjoint subsets, at least one subset is $r$-reachable. In contrast, a graph is $(r,s)$-robust if, for every pair of disjoint subsets, at least $s$ nodes have at least $r$ neighbors outside their respective subsets. Hence, $(r,s)$-robustness generalizes $r$-robustness: any $(r,s)$-robust graph with $s\geq 1$ is also $r$-robust. Conversely, if a graph is $r$-robust, it is at least $(r,1)$-robust.
While $(2F+1)$-robustness provides a sufficient condition for consensus under $F$-local Byzantine agents, $(F+1,F+1)$-robustness is both necessary and sufficient condition to tolerate $F$-total malicious agents [4]. Note that these notions represent stronger forms of network resilience compared to more traditional graph-theoretic measures such as connectivity or minimum degree [4, 49]. In particular, $r$- and $(r,s)$-robust networks are also at least $r$-connected and has minimum degree of $r$ [4, 22]. Thus, these notions of robustness enable a network to achieve the same level of resilience (i.e., tolerance to the same number of misbehaving agents) using fewer edges (cf. [49, Fig. 3]).